OpenStudy (anonymous):

Write the quadratic function in vertex form. y = x^(2) - 6x + 7

OpenStudy (jhannybean):

\[\large y= x^2 -6x +7 \]\[\large y= (x^2 -6x) +7\] complete the square. \[\large y= (x^2 -6x +9)-9+7\]Simplify.\[\large y= (x-3)^2 -2\]Vertex form : \[\large y= \color{red}a(x-\color{green}h)^2 +\color{purple}{k}\]\[\large y=\color{red}1(x-\color{green}3)^2 +(\color{purple}{-2})\]